OpenStudy (anonymous):

Please help solve \[(\frac{1}{6})^{-p}=(\frac{1}{36})^{p}\].

OpenStudy (anonymous):

i guess \(p=0\)

OpenStudy (anonymous):

\(\frac{1}{36}=(\frac{1}{6})^2\) so you have \[p=-2p\]

OpenStudy (anonymous):

I guess it doesn't say, it just says to solve the equation

OpenStudy (anonymous):

which means \(p=0\)

OpenStudy (anonymous):

i know it has to do with base changes

OpenStudy (anonymous):

i wrote the reasoning above did you understand what i wrote?

OpenStudy (anonymous):

\[\frac{1}{36}=(\frac{1}{6})^2\] so \[(\frac{1}{36})^p=(\frac{1}{6})^{-2p}\]

OpenStudy (anonymous):

i need to solve to figure out what p is equal to but I don't know how you got p=0

OpenStudy (anonymous):

the equation becomes \(-2p=p\) which is only true if \(p=0\)

OpenStudy (anonymous):

i see, thank you!